4. Find the square roots of the following numbers by the Prime Factorisation Method.  viii. 9216  

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student



9216 = 2×2×2×2×2×2×2×2×2×2×3×3×1
⇒9216=(2×2)×(2×2)×(2×2)×(2×2)×(2×2)×(3×3)
⇒9216=(2×2×2×2×2×3)×(2×2×2×2×2×3)
⇒9216=96×96
⇒√9216=96